# LangGPT: Rethinking Structured Reusable Prompt Design Framework for LLMs from the Programming Language

### v1

@@ 4 Experiments
To validate the advancement of the proposed LangGPT, we conducted experiments in two aspects: the tasks executed by LLMs and the usability of LangGPT. In our experiments, we chose two types of application scenarios: writing and role-playing. For each scenario, we selected five task-specific assistants constructed and shared by users in our community.
@@
4.1 Baselines
For comparison, we chose two baselines.
  • Instruction-only. The prompts contain only the instructions proposed for LLMs and necessary information as needed.
  • CRISPE prompt design rules Nigh (2023). The framework prospectively defines the components of the prompt from a macro perspective. It requires a complete prompt containing Capacity and Role, Insight, Statement, Personality, and Experiment.
Instruction-only prompts are very intuitive and maximize token savings. However, this approach contains too little content, often fails to convey the full range of requirements, and makes it difficult to focus on key points. CRISPE provides a relatively complete design framework rather than a number of fragmented rules. However, it does not have a clear structured design, which is not conducive to learning and the reuse of high-quality prompts. In addition, CRISPE has a strict specification of the elements to be included in the prompts, having a low generalization capability.
Figure 4: A case of a boot-licker. The responses of ChatGPT-3.5 to the user under three different prompts. Mingyuan University doesn’t really exist.
Because of the shortcomings of these methods in prompt design, we designed LangGPT. However, in addition to these intuitive advantages, we need to verify the ability of prompts to induce LLMs, which is the most essential purpose of prompt design.
4.2 Large Language Models
We selected these LLMs***All executive tasks in the experiment were completed by 18 January 2024. shown in Table 3 for evaluation.
LLM Access Scale Version
ChatGPT Schulman et al. (2023) API unk v3.5
ChatGLM Du et al. (2021) API unk v3-turbo
ChatGLM Zeng et al. (2023) Open Source 6B v3
Ernie Bot Sun et al. (2021) API unk v4.0
Gemini Anil et al. (2023) API unk Pro
Baichuan Yang et al. (2023) Open Source 7B v2-chat
Qwen Bai et al. (2023) Open Source 7B chat
Yi 01-ai (2023) Open Source 6B chat
Table 3: LLMs used in experiments. “unk” indicates that we do not know the size of those models.
4.3 Evaluation Metrics
Due to the complexity of the task scenarios, there is a lack of objective evaluation metrics. Therefore, we evaluated the ability of LLMs to execute tasks using human evaluation and LLM evaluation. To better ensure the rationality of the evaluation, we set evaluation criteria for two scenarios.
For writing tasks, we asked the evaluators to assess the effectiveness of the LLMs in performing the task in terms of three dimensions: textual continuity, formatting norms, and content richness. The ability to generate thematically coherent content is important. However, we found that LLMs in the generation task were usually able to maintain thematic consistency, so no evaluation metrics were designed for this aspect. For role-playing tasks, we similarly designed 3 dimensions of evaluation metrics: language style, characteristic relevance, and thematic coherence.
For each evaluation metric, we defined the score from zero to five to indicate different levels and indicated the description of the situation corresponding to each integer score. In addition, we allowed evaluators to score 0.5 between levels.
This evaluation framework can also guide high-performance LLMs, such as GPT-4 and Ernie Bot-4, to automatically evaluate their performance. As a result of the low consistency and of LLMs as evaluators Liu et al. (2023), we evaluate them manually.
4.4 Results
The most intuitive manifestation of the ability of a prompt to induce LLMs is the performance of LLMs in performing tasks guided by the prompt. Thus, we evaluated the performance of LLMs in two scenarios and the results are shown in Table 4.
Scenario Prompt GPT-3.5 GLM-turbo ChatGLM3 Ernie Bot Gemini Pro Baichuan2 Qwen Yi
s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3
Writing I 3.8 3.5 3.7 3.5 3.1 3.2 3.7 3.5 3.8 4.2 3.9 3.7 3.8 4.2 4.5 3.5 3.6 3.3 4.1 3.7 4.4 4.1 3.8 4.3
C 3.7 3.1 3.4 3.5 3.1 3.1 3.9 3.7 4.1 4 3.9 4 4.1 4.1 4.1 3.7 3.6 3.4 4.3 4 4.2 4.3 4 4.2
L 3.8 3.5 3.4 3.4 3.4 3.4 3.5 4 3.8 4.4 4 4.2 4.2 3.9 4.1 3.6 3.4 3.7 3.4 3.6 3.7 3.9 4.1 4.6
Role-playing I 4.2 3.5 3.4 3.8 1.6 1.4 3.4 1.4 1.7 4 0.8 2.1 4.4 2.8 3.4 3.9 1.3 1.2 3.9 2.9 2.8 3.6 2 2.4
C 4 3.7 4 4.2 2.4 2.5 4 3 3.7 4.3 2.6 2.5 4.6 4.2 4 4.1 2.5 2.4 4.3 3.2 3.4 4.2 3.2 3
L 4.3 4.3 4.3 4.2 1.8 2.1 4.1 4 4.3 4.2 2.3 2.5 4.6 4.4 4.3 4.4 3 3.2 4.2 3.7 3.7 3.8 2.8 2.8
Table 4: Results of different prompts inducing LLMs to perform tasks. Bolded fonts indicate the type of prompts that perform best in the current scenario. I, C, and L denote three kinds of prompts, i.e., Instruction-only, CRISPE, and LangGPT, respectively. $\text{s}_{1}$ to $\text{s}_{3}$ represent the three evaluation metrics for the two scenarios, respectively. For the writing scenario, $\text{s}_{1}$ is textual continuity, $\text{s}_{2}$ is formatting norms, and $\text{s}_{3}$ is content richness. For the role-playing scenario, $\text{s}_{1}$ is language style, $\text{s}_{2}$ is characteristic relevance, and $\text{s}_{3}$ is thematic coherence.
Scenario Writing Role-playing
Prompt I C L I C L
GPT-3.5 3.67 3.40 3.57 3.70 3.90 4.30
GLM-turbo 3.27 3.23 3.40 2.27 3.03 2.70
ChatGLM3 3.67 3.90 3.77 2.17 3.57 4.13
Ernie Bot 3.93 3.97 4.20 2.30 3.13 3.00
Gemini Pro 4.17 4.10 4.07 3.53 4.27 4.43
Baichuan2 3.47 3.57 3.57 2.13 3.00 3.53
Qwen 4.07 4.17 3.57 3.20 3.63 3.87
Yi 4.07 4.17 4.20 2.67 3.47 3.13
Table 5: The average of the scores of the different LLMs on the two types of Scenarios.
From the results, it can be seen that LangGPT is better at inducing LLMs to execute tasks. In addition, we noticed an interesting phenomenon. Certain LLMs with particularly strict security restrictions will refuse to answer questions or requests on topics such as bragging and crankiness. LangGPT can make these LLMs realize that it’s just a style of speaking rather than truly damaging behavior and guide them to respond. Neither of the other two baselines can do that.
@@ 4.4 Results
To evaluate the ease of use of LangGPT, we conducted a user survey in our online communityThe community has been running for more than six months and has amassed thousands of users from a wide range of industries, including manufacturing, construction, information technology, finance, and entertainment. Therefore, the objectivity of the survey results can be guaranteed.. We designed a complete questionnaire about the LangGPT experience to ensure the quality of answers. The questionnaire included a rating question on ease of use. The results of the user ratings are shown in Figure 3.
@@ 4.5 Case Study
To demonstrate the effect of LangGPT more intuitively, we filtered specific cases from our experiments. In addition to the direct effect comparison, we also tried to use LangGPT to guide LLMs to generate quality prompts.
4.5.1 Inducing LLMs
In our experiments, we guided the LLMs to play a boot-licker using three prompts, an example of which is given in Figure 4.
In this example, both the Instruction-only prompt and the CRISPE prompt-guided ChatGPT simply reply to the user’s utterances. In contrast, LangGPT-guided ChatGPT is even more bottomless in its blow-by-blow approach to the user-given subject. Additionally, it expresses compliments from a wider range of perspectives and is more in character.
Figure 5: Example of ChatGPT-3.5 to generate prompts with LangGPT. Some modules have been omitted for ease of presentation.
4.5.2 Prompt construction with LLMs
In organizing our experiments, we also tried to make LLMs construct prompts using LangGPT. An example is shown in Figure 5.
If LLMs are directly asked to generate a prompt about MBTI assessment, they might reject it. However, LangGPT can direct LLMs to generate high-quality prompts that are harmless.
@@ 5 Conclusion
In this paper, we present LangGPT, a structured and extensible framework for prompt design. LangGPT has a systematic structure similar to object-oriented programming languages and is easy to learn and reuse. Experiments demonstrate that LangGPT performs better than the baseline methods in guiding LLMs to perform tasks. We also conducted a user survey in the community built on LangGPT to verify the ease of use and reusability of LangGPT. In future work, we will further optimize the design of LangGPT and reduce the token consumption of LangGPT. In addition, support for LLMs using third-party tools and custom tools will be added.
@@
In the application of LLM, ethical disputes may arise, but the design of LangGPT and the process of writing this paper avoided possible ethical issues.

### v2

@@ 4 Experiments
To validate the advancement of the proposed LangGPT, we conducted experiments in two aspects: (1) leveraging LangGPT prompts to bootstrap LLMs to perform tasks; and (2) training LLMs on the task of generating LangGPT prompts.
4.1 Experiment Settings
Figure 2: A case of a flatterer. The responses of ChatGPT-3.5 to the user under three different prompts. Mingyuan University doesn’t really exist.
4.1.1 Large Language Models
We selected these LLMs shown in Table 4 for evaluation.
LLM Scale Version
Qwen Bai et al. (2023) 7B 1.5-chat
Deepseek DeepSeek-AI (2024) 7B chat
Llama AI@Meta (2024) 8B 3-instruct
Yi 01-ai (2023) 6B 1.5-chat
bloomz Muennighoff et al. (2022) 7B 7b1
internlm Cai et al. (2024) 7B 2-chat
Table 4: LLMs used in experiments.
(a) Average
(b) ARC-Challenge
(c) Hellaswag
(d) MMLU
(e) TruthfulQA
(f) Winogrande
(g) GSM8K
Figure 3: Results of different scales of Qwen. Each subfigure represents a different task, whereas the first subfigure represents the overall performance. ‘Instruction’ indicates that LangGPT prompts are not used while ‘LangGPT’ indicates that they are used.
4.1.2 Evaluation
There are differences in understanding of language and instructions between humans and LLMs. Thus, it is difficult and meaningless to evaluate the quality of prompts directly through metrics such as textual semantics, information richness, etc. Since the aim of prompts is to guide LLMs to perform tasks, we believe that the quality of prompts can be evaluated indirectly through the performance of LLMs. Based on this idea, we designed two methods to evaluate the quality of the proposed LangGPT. The first approach is to use LangGPT to directly instruct LLMs to perform tasks and compare the performance with that without LangGPT. In addition to this way, we constructed the LangGPT prompt generation task to train LLMs. On the one hand, we collected prompts shared by users in the LangGPT community and worked on cleaning and touching them up. On the other hand, with the help of Deepseek-V2-Chat (DeepSeek-AI, 2024), we constructed datasets for the LangGPT prompt generation using the Alpaca dataset (Taori et al., 2023) and the oaast1 dataset Köpf et al. (2023); Li et al. (2023b). The composition of the whole dataset is shown in Table 5.
Dataset Source Length Method
[…]
Table 5: The composition of LangGPT prompt generation dataset. Superscript letters indicate different subsets of the dataset, where ‘c’ denotes data from the community, and ‘o’ and ‘a’ denote data from the oaast1 and alpaca datasets, respectively.
For the evaluation of the performance of LLMs, we utilize the Open LLM Leaderboard [^2] We conducted our experiments on tasks from the first version of Open LLM Leaderboard, which has now been updated to version 2, and new experiments are in progress. (Beeching et al., 2023; Gao et al., 2021) as the main evaluation tool, which consists of ARC-challenge (Clark et al., 2018), Hellaswag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), Winogrande (Sakaguchi et al., 2019), and gsm8k (Cobbe et al., 2021).
4.2 Main Results of Performance and Analysis
The most intuitive manifestation of the quality of prompts is the performance of LLMs in performing tasks guided by them. Thus, we first evaluated the performance of LLMs in two settings to verify the quality of LangGPT prompts. The results are shown in Table 6. From the results, it can be seen that LangGPT can overall improve the performance of LLMs on executing tasks. In addition, we noticed that bloomz perform worse with the LangGPT setting. We hypothesize that this is because the LangGPT prompts are in more complex formats that may not be fully understood by low-performance models. To test this hypothesis, we conducted experiments using Qwen of different scales. The results are shown in Figure 3. From the results, it can be seen that LangGPT has a worse gain for smaller LLMs with worse overall performance. In particular, the performance is weaker than the setting without LangGPT prompts at 0.5B Qwen.
@@ 4.2 Main Results of Performance and Analysis
Model Average ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K
Qwen 53.30 45.05 76.97 60.10 53.54 65.11 19.03
Qwen 60.20 56.40 76.14 59.02 48.72 65.11 55.80
Qwen 55.38 54.35 76.80 61.30 50.17 69.77 19.86
internlm 64.00 51.28 78.31 60.55 54.09 74.66 65.13
internlm 66.00 57.00 79.76 61.67 51.93 77.19 68.46
internlm 64.29 50.94 78.38 60.38 54.79 74.51 66.72
Llama 65.99 56.83 75.80 63.84 51.65 72.06 75.74
Llama 67.13 60.32 78.34 65.31 50.03 75.22 73.54
Llama 67.94 62.29 79.26 67.41 52.51 76.48 69.67
Deepseek 56.85 49.15 77.69 50.04 47.80 69.77 46.63
Deepseek 59.28 54.35 78.38 50.77 49.53 72.53 50.11
Deepseek 60.05 55.55 79.37 52.44 46.86 76.95 49.13
Yi 65.31 54.61 77.33 63.42 52.45 71.51 72.55
Yi 66.02 59.81 78.77 61.89 50.13 71.67 73.84
Yi 66.20 61.35 77.93 64.76 49.82 74.66 68.69
bloomz 44.04 45.52 63.99 44.00 45.21 65.43 0.07
bloomz 41.93 42.74 61.73 37.01 45.78 64.33 0.00
bloomz 43.34 45.56 64.37 40.69 44.57 64.33 0.53
Table 6: Performance of LLMs at different settings. The $\bigstar$’ indicates that the original LLMs and LangGPT prompts are not used. The $\clubsuit$’ indicates that LangGPT prompts are used and the $\spadesuit$’ indicates the LLMs are fine-tuned. Bold fonts indicate results that are better than the original performance with LangGPT prompts. Underlined fonts indicate results that are better than the original performance after fine-tuning.
Moreover, we also constructed the LangGPT prompt generation task for supervised fine-tuning of LLMs in addition to directing LLMs to perform tasks. We used LoRA (Hu et al., 2021; Zheng et al., 2024) to fine-tune five open-source LLMs and test their performance. The results are shown in Table 6. From the results, the overall performance of the LLMs trained by the LangGPT prompt generation task all performed better than the original LLMs.
From the results of the two tasks above, it is clear that LangGPT prompts can help improve the performance of LLMs.
4.3 Ease of Use
@@
To evaluate the ease of use of LangGPT, we conducted a user survey in our online community. The community has been running for more than one year and has amassed thousands of users from a wide range of industries, such as manufacturing, construction, information technology, finance, entertainment, etc. Therefore, the objectivity of the survey results can be guaranteed. We designed a complete questionnaire about the LangGPT experience to ensure the quality of answers. The questionnaire included a rating question on ease of use. The results of the user ratings are shown in Figure 4.
@@ 4.4 Case Study
To demonstrate the effect of LangGPT more intuitively, we filtered specific cases from the community. We guided the LLMs to play a flatterer using LangGPT prompt and direct instruction, an example of which is given in Figure 2.
In this example, the Instruction-only prompt-guided ChatGPT showed no clear characterisation of the role and gave responses that were almost just a repetition of what the user had expressed. In contrast, LangGPT-guided ChatGPT is even more bottomless in its blow-by-blow approach to the user-given subject and expresses compliments from a wider range of perspectives.
@@ 5 Conclusion
In this paper, we present LangGPT, a dual-layer structured and extensible framework for prompt design. LangGPT has a systematic structure similar to object-oriented programming languages and is easy to learn and reuse. Experiments demonstrate that LangGPT performs better in guiding LLMs to perform tasks. We also conducted a user survey in the community built on LangGPT to verify the ease of use and reusability of LangGPT. However, experiments also show that LangGPT is currently poorly adapted to low-performance LLMs. In future work, we will further optimize the design of LangGPT, especially on low-performance LLMs. In addition, support for LLMs using third-party tools and custom tools will be added.
@@
In the application of LLM, ethical disputes may arise, but the design of LangGPT and the process of writing this paper avoided possible ethical issues. The examples given in this paper involving Mingyuan University are fictional and do not involve the evaluation or critique of any real individuals.
Limitation
The evaluation of the performance of LLMs in this paper relies on the Open LLM Leaderboard, and although it is widely adopted, the evaluation results still have some limitations.